Fruit flies have incomplete dominance in regards to eye color, where... red-eyed= RR, orange-eyed= Rr, and white-eyed= rr

Evaluate the population... 42= red, 6=orange, and 52= white

a.) What are the observed frequencies of the phenotypes? red=______ orange=______ white=________

b.) Find the proportion of R alleles (p) as well as the proportion of r alleles (q). Show work.

c.) Assuming Hardy-Weinberg equilibrium, what are the predicted frequencies of the phenotypes? red=_____ orange=______ white=________

d.) What phenomenon best describes the deviation from Hardy-Weinberg? ________________
Answer #1

2) RR the =0.44 rex = 54 -0.54 100 100 Rx 30.06 b) RA DO ADDR = 0.42 +10:06 = 045 Medecese Vr=1-0.45= 0.55 c) Expected freque

Please rate high.